# Linear Discriminant Regularized Regression

### v2

@@ B.7 Proof of Theorem 8: the reduced-rank estimator of $B$
Here $c$ is defined in 3. Lemma 20, Lemma 13 and 3 together with (3.6) ensure that $\mathcal{E}_{X}\cap\mathcal{E}_{\lambda}^{\prime}$ hold with probability at least $1-n^{-1}$. Write $\widehat{r}={\rm rank}(\widehat{B})$ and $r={\rm rank}(B)$. By the optimality of $\widehat{B}$, we have
@@
${1\over n}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}$ $\leq{2\over n}|\langle{\mathbf{Y}}-{\mathbf{X}}B,{\mathbf{X}}(\widehat{B}-B)\rangle|+\lambda(r-\widehat{r})$
$\leq 2\|\widehat{B}-B\|_{*}{1\over n}\|{\mathbf{X}}^{\top}({\mathbf{Y}}-{\mathbf{X}}B)\|_{\rm op}+\lambda(r-\widehat{r})$
$\leq 2\sqrt{r+\widehat{r}}~{}\|\widehat{B}-B\|_{F}\sqrt{\lambda/C}+\lambda(r-\widehat{r})$ by $\text{by $\mathcal{E}_{\lambda}^{\prime}$}.$
@@
$\|\widehat{B}-B\|_{F}^{2}$ $\leq{2\over c}2\sqrt{\widehat{r}+r}\|\widehat{B}-B\|_{F}\sqrt{\lambda/C}+{2\over c}\lambda(r-\widehat{r})$
$\leq{8\lambda\over c^{2}C}(\widehat{r}+r)+{1\over 2}\|\widehat{B}-B\|_{F}^{2}+{2\over c}\lambda(r-\widehat{r})$ by $2xy\leq x^{2}/2+2y^{2}$
$={1\over 2}\|\widehat{B}-B\|_{F}^{2}+{4\over c}\lambda r$ for $\text{for $C=4/c$}.$
@@
${1\over n}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}$ $\leq\sqrt{{8\over cC}\lambda(r+\widehat{r})}~{}{1\over\sqrt{n}}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}+\lambda(r-\widehat{r})$
$\leq\frac{1}{2n}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}+2\lambda r$ for $\text{for $C=4/c$}.$
@@
By (B.31), it suffices to bound from above
@@
Appendix C Technical lemmas
@@
${M}^{\top}B={M}^{\top}\Sigma_{w}^{-1}{M}~{}\Omega^{-1},$ (C.2)
$H=D_{\pi}-D_{\pi}{M}^{\top}B=\Omega^{-1},$ (C.3)
$B^{\top}\Sigma_{w}B=\Omega^{-1}{M}^{\top}\Sigma_{w}^{-1}{M}~{}\Omega^{-1}.$ (C.4)
@@
$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}\leq\pi_{\ell}\left(1\wedge\pi_{\ell}\Delta_{\ell\ell}\right),$ (C.5)
$B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}\leq\left(1\wedge{\pi_{\ell}^{2}\over\pi_{\min}}\Delta_{\ell\ell}\right),$ (C.6)
$\left|\mu_{k}^{\top}B_{\ell}\right|\leq\left(1\wedge\pi_{\ell}\sqrt{\Delta_{kk}\Delta_{\ell\ell}}\right).$ (C.7)
@@
$=\mathbf{I}_{L}-D_{\pi}^{-1}\left(D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}\right)^{-1}$ (C.8)
@@
proving (C.2). By using the above identity (C.8), we have
@@
proving (C.3). Since Lemma 1 and (C.3) imply
@@
To prove (C.5) – (C.7), pick any $k,\ell\in[L]$. By using (2.6) twice and (2.5), we find that
$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}=\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}\leq\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$ (C.9)
yielding the second bound in (C.5). The other bound in (C.5) follows by observing that
@@
For (C.6), on the one hand, similar arguments yield
@@
On the other hand, using (C.2) and (C.3) proves
@@
C.1 Concentration inequalities related with $\widehat{\Sigma}-\Sigma$
Recall that $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$ and the definition of the event $\mathcal{E}_{\pi}$ is given in (B.1).
@@ C.1 Concentration inequalities related with $\widehat{\Sigma}-\Sigma$
$\quad+\left({\|{M}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{v_{2}^{\top}\Sigma_{w}v_{2}}+{\|{M}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{v_{1}^{\top}\Sigma_{w}v_{1}}\right)\sqrt{t\over n}.$
@@
$\lesssim\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{M}^{\top}v_{1}\|_{\infty}\|{M}^{\top}v_{2}\|_{\infty}{t\over n}$ (C.11)
@@
$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}v_{2}\lesssim\|{M}^{\top}v_{1}\|_{2}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{M}^{\top}v_{1}\|_{2}\over\sqrt{m}}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n},$ (C.12)
$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}v_{2}\lesssim\|{M}^{\top}v_{2}\|_{2}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{M}^{\top}v_{2}\|_{2}\over\sqrt{m}}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}$ (C.13)
@@
For the sup-norm bound, it suffices to bound from above $|B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}|$ for any $k\in[L]$. Invoking Lemma 18 with $v_{1}=v_{2}=B_{k}$ and $t=\log(n)$ together with (C.5), (C.6) and (C.7) of Lemma 14 gives that, with probability at least $1-n^{-2}$,

### v3

@@ A.6 Proof of Theorem 8: the reduced-rank estimator of $B$
[…]
@@
[…]
@@
[…]
$={1\over 2}\|\widehat{B}-B\|_{F}^{2}+{4\over c}\lambda r$ $\text{for $C=4/c$}.$
@@
[…]
$\leq\frac{1}{2n}\|{\mathbf{X}}\widehat{B}-{\mathbf{X}}B\|_{F}^{2}+2\lambda r$ $\text{for $C=4/c$}.$
@@
By (A.31), it suffices to bound from above
@@
A.7 Proof of the equivalence between (1.1) and (4.1)
Proof.
Recall that $B^{*}=\Sigma_{w}^{-1}{M}$. We find that
so that
$\operatorname*{arg\,min}_{\ell\in[L]}~(x-\mu_{\ell})^{\top}B^{*}(B^{*\top}\Sigma_{w}B^{*})^{+}B^{*\top}(x-\mu_{\ell})$
$=\operatorname*{arg\,min}_{\ell\in[L]}~2x^{\top}\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}\mu_{\ell}+\mu_{\ell}^{\top}\Sigma_{w}^{-1}{M}({M}^{\top}\Sigma_{w}^{-1}{M})^{+}{M}^{\top}\Sigma_{w}^{-1}\mu_{\ell}.$
$\operatorname*{arg\,min}_{\ell\in[L]}~2x^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{M}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{M}^{\top}\Sigma_{w}^{-1/2}UU^{\top}\Sigma_{w}^{-1/2}{M}{\bm{e}}_{\ell}$
$=\operatorname*{arg\,min}_{\ell\in[L]}~2x^{\top}\Sigma_{w}^{-1}{M}{\bm{e}}_{\ell}+{\bm{e}}_{\ell}^{\top}{M}^{\top}\Sigma_{w}^{-1}{M}{\bm{e}}_{\ell}$
which, in view of (1.1) and (3.2), completes the proof. ∎
Appendix B Technical lemmas
@@
${M}^{\top}B={M}^{\top}\Sigma_{w}^{-1}{M}~\Omega^{-1},$ (B.2)
$H=D_{\pi}-D_{\pi}{M}^{\top}B=\Omega^{-1},$ (B.3)
$B^{\top}\Sigma_{w}B=\Omega^{-1}{M}^{\top}\Sigma_{w}^{-1}{M}~\Omega^{-1}.$ (B.4)
@@
$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}\leq\pi_{\ell}\left(1\wedge\pi_{\ell}\Delta_{\ell\ell}\right),$ (B.5)
$B_{\ell}^{\top}{M}{M}^{\top}B_{\ell}\leq\left(1\wedge{\pi_{\ell}^{2}\over\pi_{\min}}\Delta_{\ell\ell}\right),$ (B.6)
$\left|\mu_{k}^{\top}B_{\ell}\right|\leq\left(1\wedge\pi_{\ell}\sqrt{\Delta_{kk}\Delta_{\ell\ell}}\right).$ (B.7)
@@
$=\mathbf{I}_{L}-D_{\pi}^{-1}\left(D_{\pi}^{-1}+{M}^{\top}\Sigma_{w}^{-1}{M}\right)^{-1}$ (B.8)
@@
proving (B.2). By using the above identity (B.8), we have
@@
proving (B.3). Since Lemma 1 and (B.3) imply
@@
To prove (B.5) – (B.7), pick any $k,\ell\in[L]$. By using (2.6) twice and (2.5), we find that
$B_{\ell}^{\top}\Sigma_{w}B_{\ell}\leq B_{\ell}^{\top}\Sigma B_{\ell}=\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma^{-1}\mu_{\ell}\leq\pi_{\ell}^{2}\mu_{\ell}^{\top}\Sigma_{w}^{-1}\mu_{\ell}$ (B.9)
yielding the second bound in (B.5). The other bound in (B.5) follows by observing that
@@
For (B.6), on the one hand, similar arguments yield
@@
On the other hand, using (B.2) and (B.3) proves
@@
B.1 Concentration inequalities related with $\widehat{\Sigma}-\Sigma$
Recall that $\widehat{\Sigma}=n^{-1}{\mathbf{X}}^{\top}{\mathbf{X}}$ and the definition of the event $\mathcal{E}_{\pi}$ is given in (A.1).
@@ B.1 Concentration inequalities related with $\widehat{\Sigma}-\Sigma$
$\quad+\left({\|{M}^{\top}v_{1}\|_{2}\over\sqrt{L}}\sqrt{v_{2}^{\top}\Sigma_{w}v_{2}}+{\|{M}^{\top}v_{2}\|_{2}\over\sqrt{L}}\sqrt{v_{1}^{\top}\Sigma_{w}v_{1}}\right)\sqrt{t\over n}.$
@@
$\lesssim\sqrt{\sum_{k=1}^{L}(v_{1}^{\top}\mu_{k})^{2}(v_{2}^{\top}\mu_{k})^{2}\over L}\sqrt{t\over n}+\|{M}^{\top}v_{1}\|_{\infty}\|{M}^{\top}v_{2}\|_{\infty}{t\over n}$ (B.11)
@@
$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}(\widehat{\mu}_{k}-\mu_{k})\mu_{k}^{\top}v_{2}\lesssim\|{M}^{\top}v_{1}\|_{2}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{M}^{\top}v_{1}\|_{2}\over\sqrt{L}}\sqrt{tv_{2}^{\top}\Sigma_{w}v_{2}\over n},$ (B.12)
$\sum_{k=1}^{L}\widehat{\pi}_{k}v_{1}^{\top}\mu_{k}(\widehat{\mu}_{k}-\mu_{k})^{\top}v_{2}\lesssim\|{M}^{\top}v_{2}\|_{2}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}\sqrt{\max_{k}\widehat{\pi}_{k}}\lesssim{\|{M}^{\top}v_{2}\|_{2}\over\sqrt{L}}\sqrt{tv_{1}^{\top}\Sigma_{w}v_{1}\over n}$ (B.13)
@@
For the sup-norm bound, it suffices to bound from above $|B_{k}^{\top}(\widehat{\Sigma}-\Sigma)B_{k}|$ for any $k\in[L]$. Invoking Lemma 18 with $v_{1}=v_{2}=B_{k}$ and $t=\log(n)$ together with (B.5), (B.6) and (B.7) of Lemma 14 gives that, with probability at least $1-n^{-2}$,